An electromagnetic wave Here is a particular electromagnetic field in free space: (9.64) Er Eo sin at), E B2 (Eo/c) sin (kx -t cat).

(a) Show that this field can satisfy Maxwell's equations if w and k are related in a certain way.

(b) Suppose w=1010s-1 and E0=1kV/m. What is the wavelength? What is the energy density in joules per cubic meter, averaged over a large region? From this calculate the power density, the energy flow in joules per square meter per second.

(c) Show also that the electric field of associated with a spherically symmetric wave may have the dependence Ei = {Acos[k(r − ct)]}/r , where A is a constant.
Answer #1

Maxwell's equations are :

7 E E0 V B 0 V x E= t 1 OE ▽×B= μoj +2at c

a) Now let's check maxwell's equations:

ов V x E дt

ОЕ, ОЕ, ӘЕ, ДЕ, ов ДЕ, ОЕ, ј+ дл i + де k ду де дх дt ду

ов. ОEy дt дл

cos(kwt)w Encos(kawt)k

k= 3

b) when \omega = 10^{10}s^{-1}\Rightarrow \frac{2\pi}{\lambda} = \frac{10^{10}}{3 \times 10^8}\Rightarrow \lambda = \frac{6\pi}{100} = 0.19m

\text{Energy density} = \frac{1}{2} \varepsilon_{0} E^{2} + \frac{1}{2} \frac{B^{2}}{\mu_{0}} = \varepsilon_{0} E^{2} = \varepsilon_{0} E_0^{2}\sin^2 (k x+\omega t)

  Space average E 2 .   \text{Since Average of } \sin^2 \theta\text{ over large space is }\frac{1}{2}

1 B2 d(Eo sin2 (ka wt)) d(eo E2) 2 μο - Power density dt dt dt EEwsin(2(kr wt))

c)

Ei A cos k(r- ct)/r

The Wave equation is given by:

\nabla^{2} u=\frac{1}{c^{2}} \frac{\partial^{2} u}{\partial t^{2}}

\nabla^{2} E=\frac{1}{c^{2}} \frac{\partial^{2} E}{\partial t^{2}}

A Cas (-ct) ar FrSin (k-t)k AFrr-d)Gass). Cos Ckr ( (KSim(f) Cas VE = . -ArkCos (-ct AkSir-ct) t kSim Cos(-ct) Y A K Cos(k(r-